Solve by finding the basis over R for the solution space.

(A) X + 3y -3z=0

2x - 3y + z=0

3x -2y + 2z=0


(B) X + Y + Z + W=0

2X + 3Y - Z +W=0

3X + 4Y +2W=0



1
Expert's answer
2021-02-24T06:03:21-0500

"\begin{pmatrix}
 1 & 3 & -3 \\
 2 & -3 & 1 \\
 3 & -2 & 2
\end{pmatrix}
\begin{pmatrix}
x\\
y\\
z
\end{pmatrix}
=
\begin{pmatrix}
0\\
0\\
0
\end{pmatrix}"


Performing row operations on the matrix, we have


"\begin{pmatrix}
1 & 3 & -3\\
2&-3&1\\
3&-2&2
\end{pmatrix}
\to^{R_{21}(-2)}_{R_{31}(-3)}
\begin{pmatrix}
1&3&-3\\
0&-9&7\\
0&-11&11
\end{pmatrix}"


"\to R_2(\frac{-1}{9})
\begin{pmatrix}
1&3&-3\\
0&1&\frac{-7}{9}\\
0&-11&11
\end{pmatrix}
\to^{R_{32}(11)}
\begin{pmatrix}
1&3&-3\\
0&1&\frac{-7}{9}\\
0&0&\frac{22}{9}
\end{pmatrix}"


"\to R_3(\frac{9}{22})
\begin{pmatrix}
1&3&-3\\
0&1&\frac{-7}{9}\\
0&0&1
\end{pmatrix}"


"x+3y-3z=0\\
y-\frac{7}{9}z = 0\\
z=0"


Therefore x=0,y=0, z=0. This is a trivial solution hence it is not a basis

Question B

"\begin{pmatrix} 1&1&1&1\\ 2&3&-1&1\\ 3&4&0&2 \end{pmatrix} \begin{pmatrix} x\\y\\z\\w \end{pmatrix} = \begin{pmatrix} 0\\0 \\ 0 \end{pmatrix}"

The matrix of the coefficient is

"\begin{pmatrix} 1&1&1&1\\ 2&3&-1&1\\ 3&4&0&2 \end{pmatrix} \to^{R_{21}(-2)}_{R_{31}(-3)} \begin{pmatrix} 1&1&1&1\\ 0&1&-3&-1\\ 0&1&-3&-1\end{pmatrix} \to^{R_{32}(-1)} \begin{pmatrix} 1&1&1&1\\ 0&1&-3&-1\\ 0&0&0&0\end{pmatrix}"


"x+y+z+w=0\\ y-3z-w=0"

Let "z=a, w=b" where "a,b \in \mathbb{R}"

So, we have

"y-3a-b=0 \\ y = 3a+b \\ x = -y-z-w \\ x= -(3a+b) -a-b \\ x= -4a-2b\\ \therefore x = -4a-2b, y = 3a+b , z =a, w=b \\ (x,y,z,w) = a(-4,3,1,0)+b(-2,1,0,1)"

So a basis for the system is {(-4,3,1,0),(-2,1,0,1)}